Calculate the molarity and molality of an NH3 solution made up of 35.0 g of NH3 in 75.0 g of water. The density of the solution is 0.982 g/mL


1
Expert's answer
2022-02-16T07:38:02-0500

Molarity


"Mass=35+75=110g"

Volume"=110\u00d70.982=108.02ml"

Molar mass of NH"_3=17.031"


Number of moles "=\frac{35}{17.031}=2.055" moles



Molarity "=" Moles"\/" Litre

"=\frac{1000}{108.02}×2.055"

"=19.02M"


Molality "=" Moles"\/" kg of solvent

"=2.055×\frac{1000}{75}"

"=27.4" mole"\/" "kg"